OpenStudy (anonymous):

simplify the given equation . 17x-6+3x-5=x+11+4x a. 20x-11=5x+11 b. 20x+1=5x+11 c. 9x=16x show your work please. im not understanding.

pooja195 (pooja195):

Combine like terms

pooja195 (pooja195):

(17x+3x)+(−6+−5)=(x+4x)+(11)(Combine Like Terms)

OpenStudy (anonymous):

so 20x-11=5x+11?

OpenStudy (anonymous):

oh hey that was actually simple. thank you!